Net salvage value. Allen air lines must liquidate some equipment that is being replaced. The equipment originally cost $19 million of which 85% has been depreciated. The used equipment can be sold today for 6.65 million and its tax rate is 40%. What is the equipment after tax net salvage value?
Answer #1

Book Value of the equipment = 19,000,000*15% = 2,850,000

After Tax Salvage Value = (6,650,000 - 2,850,000)*(1-tax rate)  = 3,800,000*(1-40%) = 2,280,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well